Express the field D = (x2 + y2)−1(xax + yay) in cylindrical components and cylindrical

variables.


1
Expert's answer
2021-04-26T17:13:03-0400
"x = ρ \cos φ, y = ρ \sin φ\\\bold{D}=\frac{1}{\rho}(\cos φ\bold{a_x}+ \sin φ\bold{a_y})"

Then


"D_\rho=\frac{1}{\rho}(\cos φ\bold{a_x\cdot a_\rho}+ \sin φ\bold{a_y\cdot a_\rho})\\=\frac{1}{\rho}(\cos^2 φ+ \sin^2 φ)=\frac{1}{\rho}"

"D_\phi=\frac{1}{\rho}(\cos φ\bold{a_x\cdot a_\phi}+ \sin φ\bold{a_y\cdot a_\phi})\\=\frac{1}{\rho}(-\sin φ\cos φ+ \sin φ\cos φ)=0"

So,


"\bold{D}=\frac{1}{\rho}\bold{a_\rho}"